Theatre sold children ticket for $6 and adult ticket for $10. A total ticket of 80 tickets were sold for a performance bringing total $720. How many type of ticket were sold?
Adult=
Children=

Answer by mananth(16949)

Theatre sold children ticket for $6 and adult ticket for $10. A total ticket of 80 tickets were sold for a performance bringing total $720. How many type of ticket were sold?
...
children ticket $6 - x numbers
adult $10 --- 80-x numbers
..
total = $720
...
6x+10(80-x)=720
6x+800-10x=720
-4x=720-800
-4x=-80
x= 20 child tickets
60 adult tickets
...
m.ananth@hotmail.ca